OpenStudy (anonymous):

Identify which table shows a direct variation. Table 1 x y 2 2 6 12 9 18 Table 2 x y 2 14 6 42 9 63 A.Table 1 B.Table 2 C.Table 1 and Table 2 D.Neither table

OpenStudy (freckles):

x|y see for which table you have y=k*x where k is the same number over and over...

OpenStudy (freckles):

table 1: we have the following to test: see if we have y=kx where k is constant (stays the same number) x |y y=kx 2 | 2 2=k*2 6 | 12 12=k*6 9| 18 18=k*9 is the k the same value in each of the equations ?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

is it B?

OpenStudy (anonymous):

@dan815